# Coherent-state boundary conditions as the first-principles origin of background fields in QED

### v2

@@ 3.2 Transition to the Heisenberg Picture
We now focus on Eqs. (1921) and attempt to rewrite the formulation from the Schrödinger picture into the Heisenberg picture. As declared earlier, at the reference time $t_{0}$, the quantum states and operators in both pictures are identified [20, 33]. Thus
@@ 4 Furry picture
In this section, we construct the Furry picture [31, 4, 6, 9, 25] commonly employed in strong-field QED (or nonlinear QED) starting from the Heisenberg picture formulation developed in the previous section. By formulating the picture dependence of the classical background field $\mathcal{A}$ in terms of coherent states and displacement operators, we have established a consistent framework in which the spacetime dependence of $\mathcal{A}$ used in standard strong-field QED calculations is well defined and justified. In particular, while textbook formulation of QED with prescribed external fields postulate a time-dependent c-number background from the outset, here the same time dependence emerges from picture choice rather than being introduced ad hoc at the Hamiltonian level, providing a conceptual link to the standard unstable-vacuum formalism [7].
@@
In this work, laser depletion is instead addressed using scattering calculations formulated within the Schrödinger–Heisenberg framework developed in Sect. 3, with explicit applications presented in Sects. 5 and 7. For comparison, related approaches based on alternative picture choices, such as those developed by Ref. [19], will also be discussed. These frameworks allow the dynamical modification of the background field to be treated consistently and go beyond the standard Furry picture description. While we illustrate the construction using the Dirac field for concreteness, the same transformation applies equally to the EM field. Section 4.2 then shows how this general framework reproduces the familiar Volkov solutions of the Dirac field when the background field is taken to be a plane wave.
@@ 4.1 Basic structures
This formalism defines the Furry picture [31, 4, 6, 9, 25]. Use of $\hat{S}_{\mathrm{F}}$ allows the computation of scattering probabilities, for instance, nonlinear Compton scattering [31, 3, 27, 28] and nonlinear Breit–Wheeler processes [31, 27, 28, 30], as confirmed in numerous previous studies.
@@ 4.2 Volkov Solutions
When the coherent laser EM field is a plane-wave pulse, the Dirac field $\hat{\psi}_{\mathrm{F}}(x)$ admits an analytic solution in terms of the Volkov solutions [21, 31, 4, 6, 40, 20]. In this representation, the effect of the laser background is absorbed into the Dirac field itself, enabling scattering calculations to be performed with the laser influence incorporated from the outset.
@@ 5.1 Heisenberg-Euler vacuum model
where, $W[\mathcal{A}]$ is the effective action whose integrand is the HE Lagrangian density [14, 6, 31, 36, 17, 39, 5]. Within the present approach, this expression can be reformulated in terms of a coherent state. We reconsider Eq. (8) in the Schrödinger picture. Suppose that the system contains only a coherent state $|\alpha_{\mathrm{S}}(t)\rangle$ at all times. We then examine the corresponding $\alpha$-$\alpha$ transition amplitude:
@@
The goal of this subsection is not to alter the value of the HE Lagrangian, but rather to clarify its physical origin within strong-field QED when classical background fields are understood as limits of quantum-coherent EM states. From the viewpoint of coherent-field QED, the physical interpretation of the HE vacuum becomes more transparent. In the Furry picture the quantity $\langle 0_{\mathrm{F}}|\hat{V}(\infty,-\infty)|0_{\mathrm{F}}\rangle$ is described as a “vacuum–vacuum transition” in the presence of an externally prescribed background field, a somewhat formal notion in which the vacuum is said to evolve non-trivially despite containing no photons. In contrast, the coherent-field QED formulation shows that the relevant transition is, more precisely, an $\alpha$-$\alpha$ transition: a coherent photon state representing a laser pulse propagates through the vacuum while retaining its quantum state. Nothing “happens” to the vacuum itself; rather, the effective action arises from the overall phase acquired by a coherent photon state propagating through the vacuum. This interpretation resolves the long-standing conceptual ambiguity of vacuum–vacuum transitions by clarifying that the HE effective action reflects the behavior of such a coherent state, rather than any dynamical evolution of an empty vacuum.
@@ 5.2 Beyond Furry picture
An equivalent form can be found in the work of Ref. [6, 19]. The automatic inclusion of the Lorenz gauge condition is the advantage of the present method. In the present formulation, backreaction is not introduced as an independent mechanism, but appears automatically as the response of matter fields to transitions between coherent states (depletion). We note that depletion modeled via coherent-state deformations has been developed in [19]. Our treatment complements this by identifying the fixed-background approximation and its extensions as boundary-condition limits within a single first-principles framework.
@@ 6 Connection to BRST quantization
This section clarifies the relation between our coherent-state formulation and the BRST quantization of gauge theories [1, 38], ensuring the consistency of the physical-state conditions. The present formulation has been developed using the Gupta–Bleuler condition [15, 2], which provides a convenient and transparent implementation of gauge constraints in covariant QED. It is worth emphasizing, however, that this choice does not restrict the generality of the construction. In fact, the Gupta–Bleuler condition may be viewed as a special realization of the BRST formalism in the context of abelian gauge theory, where ghost degrees of freedom decouple from the physical sector (while the operator $\partial_{\mu}\hat{A}_{(+)\mathrm{H}}^{\mu}$ itself is not a Noether charge, the corresponding BRST charge provides the conserved generator that systematizes this condition into a cohomological definition of physical states).
@@
with equivalence classes taken modulo BRST-exact states. In QED, owing to the decoupling of ghost degrees of freedom, this condition reduces effectively to the Gupta–Bleuler constraint on the positive-frequency part of the gauge field [22, 26]. The coherent states considered in this work may therefore be equivalently characterized as BRST-physical states, as long as the displacement operator preserves the BRST cohomology.
@@ 7 Coherent-state boundary conditions in path-integral representation
introduced in Eq. (48) as an overlap of asymptotic in–out states in the Heisenberg picture and translate it into a path-integral representation. Here, $\alpha_{(\mathrm{in})}$ and $\alpha_{(\mathrm{out})}^{\prime}$ denote coherent states of the EM field associated with the laser modes at $t\to$-$\infty$ and $t\to+\infty$, respectively. The states $\Psi_{(\mathrm{in})}$ and $\Psi_{(\mathrm{out})}^{\prime}$ represent general multiparticle states, including electrons, positrons and photons other than the laser modes. All states are understood as asymptotic in/out states in the Heisenberg picture. From the path-integral viewpoint, fixing coherent-state boundary conditions for the EM field yields the conventional generating functional of QED in a prescribed background field as a controlled limit, while allowing $\alpha\rightarrow\alpha^{\prime}$ transitions promotes the background to a dynamical variable integrated over histories consistent with the asymptotic coherent states. This makes explicit how the standard generating functional of strong-field QED and the Heisenberg–Euler effective action arise within our framework, without double counting of photon degrees of freedom [14, 36, 7, 6].
The starting point is the observation made in the depletion analysis [6, 19]: the amplitude $S_{fi}$ corresponds to a fully quantum transition between two distinct coherent photon states, accompanied by arbitrary changes in the remaining quantum degrees of freedom. This amplitude is therefore more general than the generating functional in a fixed background field, $Z[\mathcal{A}]$, employed in conventional strong-field QED, corresponding to the fixed-background limit of coherent-field QED which implicitly assumes a fixed background field and hence $\alpha^{\prime}=\alpha$.
@@
$\hskip 0.0pt\langle\alpha_{(\mathrm{out})\mathrm{H}}^{\prime},\Psi_{(\mathrm{out})\mathrm{H}}^{\prime}|\alpha_{(\mathrm{in})\mathrm{H}},\Psi_{(\mathrm{in})\mathrm{H}}\rangle$ $=\int_{\mathcal{A}_{(-\infty)}}^{\mathcal{A}_{(+\infty)}}\mathcal{D}\mathcal{A}\int\mathcal{D}A\int\mathcal{D}\psi\int\mathcal{D}\bar{\psi}$
@@
The conventional generating functional of strong-field QED [6] is recovered as a special case. If the laser field is treated as fixed and undepleted, so that $\alpha^{\prime}=\alpha$ and the functional integration over $\mathcal{A}$ is suppressed, the above expression reduces to
$\hskip 0.0ptZ[\mathcal{A}]$ $=\int\mathcal{D}A\int\mathcal{D}\psi\int\mathcal{D}\bar{\psi}\exp\bigg[\frac{i}{\hbar c}\int d^{4}x\mathcal{L}_{\mathrm{QED}}(\psi(x),\bar{\psi}(x),A(x)+\mathcal{A}(x))\bigg]$ (55)
which is precisely the generating functional used in the Furry picture formulation and in the derivation of the Heisenberg–Euler effective action [14, 36]. From this viewpoint, $Z[\mathcal{A}]$ corresponds to a semiclassical limit of the more general coherent state transition amplitude, in which the laser field is frozen to its expectation value and quantum depletion effects are neglected.
We emphasize that the path-integral representation introduced here should be understood as a rewriting of operator-level matrix elements between physical (Gupta–Bleuler [15, 2]/BRST [1, 38, 22, 26]) states. Gauge constraints are therefore enforced prior to taking the functional integral, and the resulting expression does not represent an independent quantization procedure.
@@ 8 Conclusion
The central result of this work is that QED with prescribed classical background EM fields is not a distinct or modified theory. Rather, it is a boundary-condition limit of full QED, obtained by fixing coherent-state asymptotic boundary conditions of the EM field. This perspective differs from earlier uses of coherent states, in which the background field was treated as an externally prescribed c-number input rather than as a boundary-condition limit of full QED.
Starting from an operator formulation of QED, we have shown how the fixed background field approximation arises naturally as a limiting case of scattering between coherent states, thereby providing a precise and operator-based characterization of its conceptual status within the underlying quantum theory.
@@
Conventional generating functionals and standard formulations of QED with prescribed background fields are recovered as well-defined limiting cases corresponding to fixed coherent-state boundary conditions. Allowing for transitions between different coherent states provides a systematic extension beyond the fixed background approximation, enabling a consistent quantum-mechanical treatment of effects such as depletion and backreaction without double counting of photon degrees of freedom.
The results presented here do not introduce new properties of coherent states themselves, but instead formulate the conventional use of background fields in QED as a controlled and conceptually transparent approximation. The scope of this work is conceptual rather than computational, and we make no claim regarding improvements to perturbative techniques or explicit amplitudes beyond clarifying the operator-level structure underlying QED with prescribed background fields.

### v3

@@ 3.2 Transition to the Heisenberg Picture
We now focus on Eqs. (1921) and attempt to rewrite the formulation from the Schrödinger picture into the Heisenberg picture. As mentioned earlier, at the reference time $t_{0}$, the quantum states and operators in both pictures are identified Itzykson and Zuber (1980); Sakurai (1967). Thus
@@ 4 Furry picture
In this section, we construct the Furry picture Ritus (1985); Di Piazza et al. (2012); Fedotov et al. (2023); Furry (1951); Moortgat-Pick (2009) commonly employed in strong-field QED (or nonlinear QED) starting from the Heisenberg picture formulation developed in the previous section. By formulating the picture dependence of the classical background field $\mathcal{A}$ in terms of coherent states and displacement operators, we have established a consistent framework in which the spacetime dependence of $\mathcal{A}$ used in standard strong-field QED calculations is well defined and justified. In particular, while textbook formulation of QED with prescribed external fields postulate a time-dependent c-number background from the outset, here the same time dependence emerges from picture choice rather than being introduced ad hoc at the Hamiltonian level, providing a conceptual link to the standard unstable-vacuum formalism Fradkin et al. (1991).
@@
In this work, laser depletion is instead addressed using scattering calculations formulated within the Schrödinger–Heisenberg framework developed in Sect. 3, with explicit applications presented in Sects. 5 and 7. For comparison, related approaches based on alternative picture choices, such as those developed by Ref. Ilderton and Seipt (2018), will also be discussed. These frameworks allow the dynamical modification of the background field to be treated consistently and go beyond the standard Furry picture description. While we illustrate the construction using the Dirac field for concreteness, the same transformation applies equally to the EM field. Section 4.2 then shows how this general framework reproduces the familiar Volkov solutions of the Dirac field when the background field is taken to be a plane wave.
@@ 4.1 Basic structures
This formalism defines the Furry picture Ritus (1985); Di Piazza et al. (2012); Fedotov et al. (2023); Furry (1951); Moortgat-Pick (2009). Use of $\hat{S}_{\mathrm{F}}$ allows the computation of scattering probabilities, for instance, nonlinear Compton scattering Ritus (1985); Brown and Kibble (1964); Nikishov and Ritus (1964, 1967) and nonlinear Breit–Wheeler processes Ritus (1985); Nikishov and Ritus (1964, 1967); Reiss (1962), as confirmed in numerous previous studies.
@@ 4.2 Volkov Solutions
When the coherent laser EM field is a plane-wave pulse, the Dirac field $\hat{\psi}_{\mathrm{F}}(x)$ admits an analytic solution in terms of the Volkov solutions Joachain et al. (2011); Ritus (1985); Di Piazza et al. (2012); Fedotov et al. (2023); Wolkov (Volkov in English spelling) (1935); Itzykson and Zuber (1980). In this representation, the effect of the laser background is absorbed into the Dirac field itself, enabling scattering calculations to be performed with the laser influence incorporated from the outset.
@@ 5.1 Heisenberg-Euler vacuum model
where, $W[\mathcal{A}]$ is the effective action whose integrand is the HE Lagrangian density Greiner et al. (1985); Fedotov et al. (2023); Ritus (1985); Schwinger (1951); Heisenberg and Euler (1936); Weisskopf (1936); Dunne (2005). Within the present approach, this expression can be reformulated in terms of a coherent state. We reconsider Eq. (8) in the Schrödinger picture. Suppose that the system contains only a coherent state $|\alpha_{\mathrm{S}}(t)\rangle$ at all times. We then examine the corresponding $\alpha$-$\alpha$ transition amplitude:
@@
The goal of this subsection is not to alter the value of the HE Lagrangian, but rather to clarify its physical origin within strong-field QED when classical background fields are understood as limits of quantum-coherent EM states. From the viewpoint of coherent-field QED, the physical interpretation of the HE vacuum becomes more transparent. In the Furry picture the quantity $\langle 0_{\mathrm{F}}|\hat{V}(\infty,-\infty)|0_{\mathrm{F}}\rangle$ is described as a “vacuum–vacuum transition” in the presence of an externally prescribed background field, a somewhat formal notion in which the vacuum is said to evolve non-trivially despite containing no photons. In contrast, the coherent‑field QED formulation shows that the relevant transition is, more precisely, an $\alpha$-$\alpha$ transition: a coherent photon state representing a laser pulse propagates through the vacuum while retaining its quantum state. Nothing “happens” to the vacuum itself; rather, the effective action arises from the overall phase acquired by a coherent photon state propagating through the vacuum. This interpretation resolves the long-standing conceptual ambiguity of vacuum–vacuum transitions by clarifying that the HE effective action reflects the behavior of such a coherent state, rather than any dynamical evolution of an empty vacuum.
@@ 5.2 Beyond Furry picture
An equivalent form can be found in the work of Ref. Fedotov et al. (2023); Ilderton and Seipt (2018). The automatic inclusion of the Lorenz gauge condition is the advantage of the present method. In the present formulation, backreaction is not introduced as an independent mechanism, but appears automatically as the response of matter fields to transitions between coherent states (depletion). We note that depletion modeled via coherent-state deformations has been developed in Ilderton and Seipt (2018). Our treatment complements this by identifying the fixed background-field approximation and its extensions as boundary-condition limits within a single first-principles framework.
@@ 6 Connection to BRST quantization
This section clarifies the relation between our coherent-state formulation and the BRST quantization of gauge theories Becchi et al. (1976); Tyutin (1975), ensuring the consistency of the physical-state conditions. The present formulation has been developed using the Gupta–Bleuler condition Gupta (1950); Bleuler (1950), which provides a convenient and transparent implementation of gauge constraints in covariant QED. It is worth emphasizing, however, that this choice does not restrict the generality of the construction. In fact, the Gupta–Bleuler condition may be viewed as a special realization of the BRST formalism in the context of abelian gauge theory, where ghost degrees of freedom decouple from the physical sector (while the operator $\partial_{\mu}\hat{A}_{(+)\mathrm{H}}^{\mu}$ itself is not a Noether charge, the corresponding BRST charge provides the conserved generator that systematizes this condition into a cohomological definition of physical states).
@@
with equivalence classes taken modulo BRST-exact states. In QED, owing to the decoupling of ghost degrees of freedom, this condition reduces effectively to the Gupta–Bleuler constraint on the positive-frequency part of the gauge field Kugo and Ojima (1979); Nakanishi and Ojima (1990). The coherent states considered in this work may therefore be equivalently characterized as BRST-physical states, as long as the displacement operator preserves the BRST cohomology.
@@ 7 Coherent-state boundary conditions in path-integral representation
introduced in Eq. (48) as an overlap of asymptotic in–out states in the Heisenberg picture and translate it into a path-integral representation. Here, $\alpha_{(\mathrm{in})}$ and $\alpha_{(\mathrm{out})}^{\prime}$ denote coherent states of the EM field associated with the laser modes at $t\to$-$\infty$ and $t\to+\infty$, respectively. The states $\Psi_{(\mathrm{in})}$ and $\Psi_{(\mathrm{out})}^{\prime}$ represent general multiparticle states, including electrons, positrons and photons other than the laser modes. All states are understood as asymptotic in/out states in the Heisenberg picture. From the path-integral viewpoint, fixing coherent-state boundary conditions for the EM field yields the conventional generating functional of QED in a prescribed background field as a controlled limit, while allowing $\alpha\rightarrow\alpha^{\prime}$ transitions promotes the background to a dynamical variable integrated over histories consistent with the asymptotic coherent states. This makes explicit how the standard generating functional of strong-field QED and the Heisenberg–Euler effective action arise within our framework, without double counting of photon degrees of freedom Greiner et al. (1985); Schwinger (1951); Fradkin et al. (1991); Fedotov et al. (2023).
The starting point is the observation made in the depletion analysis Fedotov et al. (2023); Ilderton and Seipt (2018): the amplitude $S_{fi}$ corresponds to a fully quantum transition between two distinct coherent photon states, accompanied by arbitrary changes in the remaining quantum degrees of freedom. This amplitude is therefore more general than the generating functional in a fixed background field, $Z[\mathcal{A}]$, employed in conventional strong-field QED, corresponding to the fixed background-field limit of coherent-field QED which implicitly assumes a fixed background-field and hence $\alpha^{\prime}=\alpha$.
@@
$\langle\alpha_{(\mathrm{out})\mathrm{H}}^{\prime},\Psi_{(\mathrm{out})\mathrm{H}}^{\prime}|\alpha_{(\mathrm{in})\mathrm{H}},\Psi_{(\mathrm{in})\mathrm{H}}\rangle$ $=\int_{\mathcal{A}_{(-\infty)}}^{\mathcal{A}_{(+\infty)}}\mathcal{D}\mathcal{A}\int\mathcal{D}A\int\mathcal{D}\psi\int\mathcal{D}\bar{\psi}$
@@
The conventional generating functional of strong-field QED Fedotov et al. (2023) is recovered as a special case. If the laser field is treated as fixed and undepleted, so that $\alpha^{\prime}=\alpha$ and the functional integration over $\mathcal{A}$ is suppressed, the above expression reduces to
$Z[\mathcal{A}]$ $=\int\mathcal{D}A\int\mathcal{D}\psi\int\mathcal{D}\bar{\psi}\exp\bigg[\frac{i}{\hbar c}\int d^{4}x\mathcal{L}_{\mathrm{QED}}(\psi(x),\bar{\psi}(x),A(x)+\mathcal{A}(x))\bigg]$ (55)
which is precisely the generating functional used in the Furry picture formulation and in the derivation of the Heisenberg–Euler effective action Greiner et al. (1985); Schwinger (1951). From this viewpoint, $Z[\mathcal{A}]$ corresponds to a semiclassical limit of the more general coherent state transition amplitude, in which the laser field is frozen to its expectation value and quantum depletion effects are neglected.
We emphasize that the path-integral representation introduced here should be understood as a rewriting of operator-level matrix elements between physical (Gupta–Bleuler Gupta (1950); Bleuler (1950)/BRST Becchi et al. (1976); Tyutin (1975); Kugo and Ojima (1979); Nakanishi and Ojima (1990)) states. Gauge constraints are therefore enforced prior to taking the functional integral, and the resulting expression does not represent an independent quantization procedure.
@@ 8 Conclusion
The central result of this work is that QED with prescribed classical background EM fields is not a distinct or modified theory. Rather, it is a boundary-condition limit of full QED, obtained by fixing coherent-state asymptotic boundary conditions of the EM field. This perspective differs from earlier uses of coherent states, in which the background field was treated as an externally prescribed c-number input rather than as a boundary‑condition limit of full QED.
Starting from an operator formulation of QED, we have shown how the fixed background-field approximation arises naturally as a limiting case of scattering between coherent states, thereby providing a precise and operator-based characterization of its conceptual status within the underlying quantum theory.
@@
Conventional generating functionals and standard formulations of QED with prescribed background fields are recovered as well-defined limiting cases corresponding to fixed coherent-state boundary conditions. Allowing for transitions between different coherent states provides a systematic extension beyond the fixed background-field approximation, enabling a consistent quantum-mechanical treatment of effects such as depletion and backreaction without double counting of photon degrees of freedom.
The results presented here do not introduce new properties of coherent states themselves, but instead formulate the conventional use of background fields in QED as a controlled and conceptually transparent approximation. The scope of this work is conceptual rather than computational, and we make no claim regarding improvements to perturbative techniques or explicit amplitudes beyond clarifying the operator-level structure underlying QED with prescribed background fields. At the conceptual level, the coherent-state transitions underlying the fixed background-field approximation are intrinsically nonperturbative in nature. In this sense, the fixed background description of QED appears as a genuinely nonperturbative boundary-condition limit of the full theory. The coherent-state specification of the asymptotic EM field conditions encodes vacuum persistence, depletion and pair creation in a way that does not rely on any perturbative expansion, but instead on the global structure of the Hilbert space.